# Measuring the Impact of Missingness in Traffic Stop Data

### v1

@@ 3.3 Notable dCMR trends across the SOPP
We reviewed the dCMRs for all datasets and selected the six plots in Figure 4 as representative of common MNAR trends. The problematic relationships between dCMR and date (by week) include piecewise constant; changes in variance; and non-monotonic trends across time.
@@
In Figure 5 we demonstrate the usefulness of the maximal correlation by selecting the five plots corresponding to the minimum maximal correlation, the maximal correlation quartiles (25th, 50th, and 75th percentiles), and the maximum maximal correlation, providing a high level summary of the extent of dependence between missingness and date. One might argue that we cherry-picked the datasets in Figure 4. However, the maximal correlation values in Figure 4 are around the $50^{th}$ and $75^{th}$ quantiles of the maximal correlation distribution (seen in Figure 5), indicating that the datasets described in Figure 4 are not extreme cases compared to the overall trends we see across all of the datasets.
@@ 3.4.1 Temporal trends
First, we examine how missingness correlates with date in Figure 6. Although the scale of the y-axis is small for many of these plots, we clearly see non-random missingness trends over time. Three datasets (California, Colorado, and Washington) have increasing missingness over time; two (Chicago and New Jersey) have decreasing; and three (Maryland, Ohio, and Wisconsin) roughly equal trends over time.
Figure 6: The dCMR of eight datasets conditional on date (binned into weeks). The maximal correlation for each dataset is included in parantheses.
When considered across time (binned by hour), the dCMRs also exhibit dependence with the conditioning variable in Figure 7. In four datasets (Chicago, New Jersey, Ohio, and Wisconsin), missingness peaks in the early mornings; in two datasets (Maryland and Washington state), missingness has two local peaks – one in the early morning, and one later in the day.
@@ 3.4.2 Spatial trends
When we condition on latitude and longitude, we can plot each bin (6-digit geohash) with the $x$-axis as longitude and the $y$-axis as latitude. Then, the dCMR can be color-coded. For the three state-level datasets in Figure 8, the dCMR plot on latitude and longitude naturally recovers state highways. The maximal correlation here is an average of the maximal correlation between latitude and dCMR and between longitude and dCMR. The maximal correlation between location and missingness in state-wide datasets is low, with only Wisconsin state having maximal correlation above 0.1. Near city centers like Minneapolis towards the west and Milwaukee in the east, the dCMR is slightly higher than its surroundings. Meanwhile, in Chicago, one can clearly see the higher levels of missingness in the city center and lower missingness in the suburbs.
@@ 4 Assessing sensitivity of missingness
Using the SOPP, we analyze several datasets to broadly explore differential missingness. As a way to investigate the impact of the missing race data, we assign observations whose race is not measured to either the Black or white groups of individuals, randomly. Subsequently, we apply two published methods to the augmented dataset in order to investigate how missingness can impact racial bias estimates. Although both published studies consider race within the framework of missing potential outcomes, neither study delves into the impact of missing data in considering the reporting of a race effect. Our sensitivity analysis furthers the literature on understanding racial discrimination in traffic policing.
@@ 4.1 The outcome test
An analysis of over 20 million traffic stops applies the outcome test (originally described in Becker, (1993) applied to traffic data in Pierson et al., (2020, Figure 3)) to compare the rate of searches to the rate of successful searches in order to determine if the police officers use a lower standard for searching drivers of color (race effect). The researchers find evidence of both racial profiling and preferential searches for minority race drivers. Although the discussion in Pierson et al., (2020) focuses on standardizing and improving the quality of traffic stop data collection, it does not mention how missing data plays a role in the analyses.
Applied to the traffic data, the outcome test (Pierson et al.,, 2020, Figure 3) compares “hit rates” (point estimates of the probability of recovering contraband given that a driver is searched) between Black and white drivers. The idea is that if searches of minority drivers reveal contraband at lower rates than searches of white drivers, the outcome test suggests that officers are searching minority drivers on the basis of less evidence and with lower success, i.e., they are potentially racially biased.
The outcome test is not designed to identify causal mechanisms for differential hit rates; instead, it is a measure of the race effect. For example, a lower hit rate could be due to racial bias of the officers or because Black drivers are carrying around less contraband than white drivers. However, the latter reason would require that (1) the police be colorblind and (2) the police are not using any other correlated indicators related to contraband (e.g., the age of the vehicle). Our work does not delve into the shortcomings of the outcome test (e.g., inability to identify causal indicators), which has been studied in detail in prior work (Engel,, 2008). Instead, we offer a sensitivity analysis that assesses the impact of missing data race information. We offer our sensitivity analysis to complement analyses such as Becker, (1993), Persico and Todd, (2006), and Pierson et al., (2020) who use the outcome test in their own work.
@@ 4.1.1 Sensitivity analysis of the outcome test
Throughout our sensitivity analysis, we use disparity as a measure of racial discrimination. Equation (3) provides the generic form of disparity where $p$ represents the probability of some police action (e.g., being searched). Equations (4) and (5) are specific to the outcome test (here, Section 4.1.1) and bounds on the ATE (see Section 4.2.1), respectively. It is important to note, however, that $\text{disparity}_{\text{outcome}}<0$ indicates discrimination against Black drivers because if Black drivers have lower rates of contraband found than white drivers, they are being stopped unfairly. On the other hand, $\text{disparity}_{\text{ATE}}>0$ indicates discrimination against Black drivers because the traffic stop “as if Black” is being searched more often than the traffic stop “as if white.”
@@
[…]
The first step of the sensitivity analysis is to calculate the ignore NA disparity: point estimate of the probability of recovering contraband given that a driver is searched for Black drivers minus white drivers (without considering any observations with NA race). A negative disparity signals discrimination against Black drivers because it indicates that Black drivers are more often searched without having contraband found.
@@
{adjustwidth}
-2.5 cm-2.5 cm[!htb]Outcome test disparity estimates. $d=p_{black}-p_{white}>0$ ($+$) indicating discrimination against white drivers; $d=p_{black}-p_{white}<0$ (--) indicating discrimination against Black drivers. Dataset Total number of counties/ departments Counties/ departments with missingness Sign of ignore NA disparity estimate Imputed county-level disparity estimate (--) ($+$) Counties that could switch signs Counties that don’t switch signs (--) to ($+$) ($+$) to (--) Remains (--) Remains ($+$) OH 89 87 10/87 77/87 5 66 5 11 CO 66 32 25/32 7/32 0 2 25 5 WI 72 49 25/49 24/49 3 2 22 22 WA 39 31 19/31 12/31 3 3 16 9 MD 223 60 34/60 26/60 0 2 34 24
@@ 4.2 Bounds on the average treatment effect (ATE)
Knox et al., (2020) conceptualize race as a causal variable (missing potential outcomes) and the “investigation of racial bias in policing as an inherently causal inquiry.” Using causal inference, they derive strict bounds for the Average Treatment Effect (ATE) (the extent to which drivers of color face greater risk of force than white drivers because of their race), given a traffic stop was made. In particular, the ATE measures the probability of being searched if a person was Black and stopped as if they were Black minus the probability of being searched if a person was white and stopped as if they were white, given that a stop was made, see Equation 5. They do not discuss missing data or its impact on the analyses. However, they do require the “Mandatory Reporting” assumption (that all encounters are recorded) in their derivations (an assumption which is clearly violated if some data are missing).
Knox et al., (2020) describe the missing potential outcomes seen in Figure 3 as the triangle between $X_{2}$, $M$, and $Y$. We extend their work to include missing data ($R_{X_{2}}$, $R_{X}$, and $R_{Y}$). Because it is not possible to recover the relationship between outcome and treatment in the missing data / missing potential outcomes scenario (Moreno-Betancur et al.,, 2018), we provide simulated bounds to assess the impact of $R_{X_{2}}$, missing data.
Knox et al., (2020) derive nonparametric bounds of ATE (given a traffic stop was made) on being searched, as a function of $\rho$, the proportion of racially discriminatory stops. Note that $\rho$ is unknowable, and the Knox et al., (2020) results are provided over a range of possible $\rho$ values. As such, they are able to measure the missing potential outcomes information (in the complete data case), but they do not evaluate the impact of missing data on the missing potential outcomes analysis.
@@ 4.2.1 Sensitivity analysis of the bounds on average treatment effect (ATE)
To address possible non-MCAR missingness in race on ATE, we perform another sensitivity analysis where we calculate the ATE bounds when NA values have been ignored (equivalent to Knox et al., (2020) which are based on bounds by Horowitz and Manski, (2000)) and compare with the bounds when NA values have been allocated to the Black and white groups. Consider Colorado and Washington highway patrols, as seen in Figure 10. Following Knox et al., (2020), by ignoring the NA race values, we calculate the naïve disparity and also bounds on the disparity for each of three different values of $\rho$ (the proportion of racially discriminatory stops): 0.25, 0.5, 0.75. Considering a higher proportion of racially discriminatory stops naturally leads to a higher range of values of disparity.
@@
For computational reasons, we were unable to run the ATE disparity estimates for every possible allocation of NA race drivers. However, along with the sensitivity analysis showing how disparity changes depending on the proportion of NA race drivers assigned to white, we also analyzed how the disparity would change under the most extreme cases. In particular, we added two “extreme” measurements to each sensitivity plot where the NA race drivers who were searched were all allocated to the Black driver group and the NA race drivers who were not searched were all allocated to the white driver group.
Note that the “extreme” cases (Figure 10 and Appendix D) seem to violate the strict bounds of Knox et al., (2020). However, the bounds from Knox et al., (2020) are only strict if there is no missingness (or if the missingness is MCAR) and are derived given the DAG of Knox et al., (2020) which is the red subset of the DAG in Figure 3. In Figure 3 and Section 3 we justified arrows from the variables to the missingness indicators which motivate our sensitivity analysis. And the sensitivity analysis demonstrates that the strict bounds of Knox et al., (2020) are not strict if the NA race observations are considered.
@@ 5 Discussion
Our work demonstrates how missing data information can be used to understand missingness trends and the sensitivity of racial disparity outcomes to missing data. We clarify the distinction between missing data and missing potential outcomes and reason that understanding both is important. Our sensitivity analysis is applied to two metrics of racial discrimination: the outcome test (Pierson et al.,, 2020) and bounds on the ATE (Knox et al.,, 2020). There are many other metrics which have been used to determine racial bias in traffic stops and could be assessed using a similar sensitivity analysis, for example, the KPT test (Knowles et al.,, 2001), the Threshold test (Simoiu et al.,, 2017; Pierson et al.,, 2018, 2020), Veil of Darkness (Grogger and Ridgeway,, 2006), Internal Benchmarking (Ridgeway and MacDonald,, 2010), or the Hit Rates test (Persico and Todd,, 2006).
Additionally, the methods provided here are applied exclusively to traffic stop data. However, the ideas could be expanded to data of many different types. As mentioned in Section 2, we see the importance of addressing missingness in healthcare (Kaambwa et al.,, 2012; Ward et al.,, 2020; Hunt et al.,, 2021; Zhang et al.,, 2023), natural disasters (Jones et al.,, 2022), psychology (Gomer and Yuan,, 2023), agriculture (Robbins et al.,, 2013), epidemiology (Allotey and Harel,, 2019; Balzer et al.,, 2020; Tsiampalis and Panagiotakos,, 2020), and crime (Edwards et al.,, 2017; Blasco,, 2021; Stockton et al.,, 2024). A sensitivity analysis using the steps described in Figure 1 could be valuable across these and many other disciplines in which missing data convolutes the resulting analysis conclusions.
@@ Acknowledgments
S.K. was supported in part by grants from the Pomona College SURP program, the Kenneth Cooke Summer Research Fellowship, and NSF Conference grant 2328509. Additionally, we appreciate the ideas put forth by Ghassan Sarkis and the following students, who worked on aspects of the project: Oliver Chang, Emma Godfrey, Will Gray, Ethan Ong, Kyle Torres, Arm Wonghirundacha, and Ivy Yuan.
@@
S.K. conducted the computational study and wrote the related aspects of the manuscript. A.L. introduced the missing data metrics and wrote the related aspects of the manuscript. J.H. conceived of the investigation, supported the computational work, and wrote sections of the manuscript.

### v2

@@ 3.3 Notable dCMR trends across the SOPP
We reviewed the dCMRs for all datasets and selected the six plots in Figure 4 as representative of common trends within the SOPP datasets (all of which seem to be be not MCAR). The problematic relationships between dCMR and date (by week) include piecewise constant; changes in variance; and non-monotonic trends across time. The information within the data itself does not provide reasons for the patterns; they could be due to procedural changes, changes in how the data were collected, or changes in missingness for particular subsets of variables. If the research goal is to truly understand the exact missingness patterns, we recommend a further dive into the CMR plots for individual variables. Note, however, that our work has found no strong trends between the additional variables and the variables of interest in our work (i.e., race and outcome measures like searched and contraband found). The dCMR (missingness on the y-axis) for each of the plots is on a different scale. We let the scales be free in order to focus on the trends in the missingness patterns.
@@
In Figure 5 we demonstrate the usefulness of the maximal correlation by selecting the five plots corresponding to the minimum maximal correlation, the maximal correlation quartiles (25th, 50th, and 75th percentiles), and the maximum maximal correlation, providing a high level summary of the extent of dependence between missingness and date. One might argue that we cherry-picked the datasets in Figure 4. However, the maximal correlation values in Figure 4 are around the $50^{th}$ and $75^{th}$ quantiles of the maximal correlation distribution (seen in Figure 5), indicating that the datasets described in Figure 4 are not extreme cases compared to the overall trends we see across all of the datasets. Note that the dCMR (missingness on the y-axis) for each of the plots is on a different scale. We let the scales be free in order to focus on the trends in the missingness patterns.
@@ 3.4.1 Temporal trends
First, we examine how missingness correlates with date in Figure 6. Although the scale of the y-axis is small for many of these plots, we clearly see non-random missingness trends over time. Three datasets (California, Colorado, and Washington) have increasing missingness over time; two (Chicago and New Jersey) have decreasing; and three (Maryland, Ohio, and Wisconsin) have trends that are roughly equal over time, with each state having a handful of years that are clearly outliers with respect to the missingness. As with Figure 4, we cannot know the reason for the missingness patterns. A deeper dive into the CMR plots for individual variables could shed light into some of the relationships, however, we have found that none of the variables are strongly associated with the missingness trends connecting race and outcome variables (like search and contraband found).
Figure 6: The dCMR of eight datasets conditional on date (binned into weeks). The maximal correlation for each dataset is included in parentheses.
When considered across time (binned by hour), the dCMRs also exhibit dependence with the conditioning variable in Figure 7. In four datasets (Chicago, New Jersey, Ohio, and Wisconsin), missingness peaks in the early mornings; in two datasets (Maryland and Washington state), missingness has two local peaks – one in the early morning, and one later in the day. However, we point out that the range of dCMR missingness is quite small, and the peaks in missingness may have to do with external factors.
@@ 3.4.2 Spatial trends
When we condition on latitude and longitude, we can plot each bin (6-digit geohash) with the $x$-axis as longitude and the $y$-axis as latitude. Then, the dCMR can be color-coded. For the three state-level datasets in Figure 8, the dCMR plot on latitude and longitude naturally recovers state highways (although it is not a perfect representation due to some highway patrols working primarily on highways (WA) and some patrolling side streets as well (OH)). The maximal correlation here is an average of the maximal correlation between latitude and dCMR and between longitude and dCMR. The maximal correlation between location and missingness in state-wide datasets is low, with only Wisconsin state having maximal correlation above 0.1. Near city centers like Minneapolis towards the west and Milwaukee in the east, the dCMR is slightly higher than its surroundings. Meanwhile, in Chicago, one can clearly see the higher levels of missingness in the city center and lower missingness in the suburbs.
@@ 4 Assessing sensitivity of missingness
Using the SOPP, we analyze several datasets to broadly explore differential missingness. As a way to investigate the impact of the missing race data, we assign observations whose race is not measured to either the Black or white groups of individuals, randomly. Subsequently, we apply two published methods to the augmented dataset in order to investigate how missingness can impact racial bias estimates. Although both published studies consider race within the framework of missing potential outcomes, neither study delves into the impact of missing data in considering the reporting of a race effect. Our sensitivity analysis furthers the literature on understanding racial discrimination in traffic policing. Our work has similarities to meyer2024 who use a sensitivity analysis on a new test, the Overlapping Condition Test.
@@ 4.1 The outcome test
An analysis of over 20 million traffic stops applies the outcome test (originally described in becker1993nobel applied to traffic data in pierson2020large) to compare the rate of searches to the rate of successful searches in order to determine if the police officers use a lower standard for searching drivers of color (race effect). The researchers find evidence of both racial profiling and preferential searches for minority race drivers. Although the discussion in pierson2020large focuses on standardizing and improving the quality of traffic stop data collection, it does not mention how missing data plays a role in the analyses.
Applied to the traffic data, the outcome test (pierson2020large, Figure 3) compares “hit rates” (point estimates of the probability of recovering contraband given that a driver is searched) between Black and white drivers. The idea is that if searches of minority drivers reveal contraband at lower rates than searches of white drivers, the outcome test suggests that officers are searching minority drivers on the basis of less evidence and with lower success, i.e., they are potentially racially biased.
The outcome test is not designed to identify causal mechanisms for differential hit rates; instead, it is a measure of the race effect. For example, a lower hit rate could be due to racial bias of the officers or because Black drivers are carrying around less contraband than white drivers. However, the latter reason would require that (1) the police be colorblind and (2) the police are not using any other correlated indicators related to contraband (e.g., the age of the vehicle). Our work does not delve into the shortcomings of the outcome test (e.g., inability to identify causal indicators), which have been studied in detail in prior work (engel2008critique). Instead, we offer a sensitivity analysis that assesses the impact of missing data race information. We offer our sensitivity analysis to complement analyses such as becker1993nobel, persico2006generalising, pierson2020large, and gaebler2025 who use the outcome test in their own work.
@@ 4.1.1 Sensitivity analysis of the outcome test
Throughout our sensitivity analysis, we use disparity as a measure of racial discrimination. Equation (3) provides the generic form of disparity where $p$ represents the probability of some police action (e.g., being searched). Equations (4) and (5) are specific to the outcome test (here, Section 4.1.1) and bounds on the ATE (see Section 4.2.1), respectively. It is important to note, however, that $\text{disparity}_{\text{outcome}}<0$ indicates discrimination against Black drivers because if Black drivers have lower rates of contraband found than white drivers, they are being searched unfairly. On the other hand, $\text{disparity}_{\text{ATE}}>0$ indicates discrimination against Black drivers because the traffic stop “as if Black” is being searched more often than the traffic stop “as if white.”
@@
$\text{disparity}_{\text{outcome}}$ $=$ $P(\text{contraband }|\text{ stop, search},\text{Black})-P(\text{contraband }|\text{ stop, search},\text{white})$ (4)
@@
The first step of the sensitivity analysis is to calculate the ignore NA disparity: point estimate of the probability of recovering contraband given that a driver is searched for Black drivers minus white drivers (without considering any observations with NA race).
@@
[!htb]Outcome test disparity estimates. $d=p_{Black}-p_{white}>0$ (++) indicating discrimination against white drivers; $d=p_{Black}-p_{white}<0$ (-) indicating discrimination against Black drivers. Dataset Total number of counties/ departments Counties/ departments with missingness Sign of ignore NA disparity estimate Imputed county-level disparity estimate (-) (++) Counties that could switch signs Counties that don’t switch signs (-) to (++) (++) to (-) Remains (-) Remains (++) OH 89 87 10/87 77/87 5 66 5 11 CO 66 32 25/32 7/32 0 2 25 5 WI 72 49 25/49 24/49 3 2 22 22 WA 39 31 19/31 12/31 3 3 16 9 MD 223 60 34/60 26/60 0 2 34 24
@@ 4.2 Bounds on the average treatment effect (ATE)
knox2020administrative conceptualize race as a causal variable (missing potential outcomes) and the “investigation of racial bias in policing as an inherently causal inquiry.” Using causal inference, they derive strict bounds for the Average Treatment Effect (ATE) (the extent to which drivers of color face greater risk of force than white drivers because of their race), given a traffic stop was made. In particular, the ATE measures the probability of being searched if a person was Black and stopped as if they were Black minus the probability of being searched if a person was white and stopped as if they were white, given that a stop was made, see Equation 5. They do not discuss missing data or its impact on the analyses. However, they do require the “Mandatory Reporting” assumption (that all encounters are recorded) in their derivations (an assumption which is clearly violated if some data are missing).
knox2020administrative describe the missing potential outcomes seen in Figure 3 as the triangle between $X_{2}$, $M$, and $Y$. We extend their work to include missing data ($R_{X_{2}}$, $R_{X}$, and $R_{Y}$). Because it is not possible to recover the relationship between outcome and treatment in the missing data / missing potential outcomes scenario (moreno2018canonical), we provide simulated bounds to assess the impact of $R_{X_{2}}$, missing data.
knox2020administrative derive nonparametric bounds of ATE (given a traffic stop was made) on being searched, as a function of $\rho$, the proportion of racially discriminatory stops. Note that $\rho$ is unknowable, and the knox2020administrative results are provided over a range of possible $\rho$ values. As such, they are able to measure the missing potential outcomes information (in the complete data case), but they do not evaluate the impact of missing data on the missing potential outcomes analysis.
@@ 4.2.1 Sensitivity analysis of the bounds on average treatment effect (ATE)
To address possible non-MCAR missingness in race on ATE, we perform another sensitivity analysis where we calculate the ATE bounds when NA values have been ignored (equivalent to knox2020administrative which are based on bounds by horowitz2000nonparametric) and compare with the bounds when NA values have been allocated to the Black and white groups. Consider Colorado and Washington highway patrols, as seen in Figure 10. Following knox2020administrative, by ignoring the NA race values, we calculate the naïve disparity and also bounds on the disparity for each of three different values of $\rho$ (the proportion of racially discriminatory stops): 0.25, 0.5, 0.75. For each of three different values of $\rho$ we calculate the bounds on the ATE, while ignoring the NA values (just as knox2020administrative). Those bounds are represented by the dashed vertical lines and the corresponding ribbon plots. The solid vertical lines are the Knox et al. ATE bounds at each of a different value for the proportion NA assigned to white. The red dots represent the estimated ATE (naïve estimate) at each value for the proportion NA assigned to white. As expected, a higher proportion of racially discriminatory stops naturally leads to a higher range of values of disparity.
@@
Note that the “extreme” cases (Figure 10 and Appendix D) seem to violate the strict bounds of knox2020administrative. However, the bounds from knox2020administrative are only strict if there is no missingness (or if the missingness is MCAR) and are derived given the DAG of knox2020administrative which is the red subset of the DAG in Figure 3. In Figure 3 and Section 3 we justified arrows from the variables to the missingness indicators which motivate our sensitivity analysis. And the sensitivity analysis demonstrates that the strict bounds of knox2020administrative are not strict if the NA race observations are considered.
@@ 5 Discussion
Our work demonstrates how missing data information can be used to understand missingness trends and the sensitivity of racial disparity outcomes to missing data. We clarify the distinction between missing data and missing potential outcomes and reason that understanding both is important. Our sensitivity analysis is applied to two metrics of racial discrimination: the outcome test (pierson2020large) and bounds on the ATE (knox2020administrative). There are many other metrics which have been used to determine racial bias in traffic stops and could be assessed using a similar sensitivity analysis, for example, the KPT test (knowles2001), the Threshold test (simoiu2017; pierson2018; pierson2020large), Internal Benchmarking (ridgeway2010methods), or the Hit Rates test (persico2006generalising).
Note that there are methods, such as the Veil of Darkness (grogger2006testing), which have been able to directly address the problem of data reporting rates that vary by race through modeling. In particular, their model uses ratios of probabilities (of data reporting) that can vary by race but do not vary given race and time. Not every model, however, can easily integrate probabilities of reporting, and we believe that our method of running a sensitivity analysis can benefit many models which measure racial disparities in policing.
Additionally, the methods provided here are applied exclusively to traffic stop data. However, the ideas could be expanded to data of many different types. As mentioned in Section 2, we see the importance of addressing missingness in healthcare (kaambwa2012; ward2020; hunt2021; zhang2023machine), natural disasters (jones2022human), psychology (gomer2023realistic), agriculture (robbins2013imputation), epidemiology (allotey2019multiple; balzer2020; tsiampalis2020missing), and crime (edwards2017effect; blasco2021missing; stockton2024now). A sensitivity analysis using the steps described in Figure 1 could be valuable across these and many other disciplines in which missing data convolutes the resulting analysis conclusions.
@@ Acknowledgments
S.K. was supported in part by grants from the Pomona College SURP program, the Kenneth Cooke Summer Research Fellowship, and NSF Conference grant 2328509. Amber Lee was important to much of the work, and we appreciate her early contributions. Additionally, we appreciate the ideas put forth by Ghassan Sarkis and the following students, who worked on aspects of the project: Oliver Chang, Emma Godfrey, Will Gray, Ethan Ong, Kyle Torres, Arm Wonghirundacha, and Ivy Yuan.
@@
S.K. conducted the computational study and wrote the related aspects of the manuscript. J.H. conceived of the investigation, supported the computational work, and wrote sections of the manuscript.